OpenStudy (anonymous):

-14x^3/x^3 - 5x^4 A.-14/5x - 1; where x = 1/5, 0 B. -14x/1 - 5x; where x = 1/5 C. 1 - 5x/-14x; where x = 0 D. -14/1 - 5x; where x = 1/5, 0 All the equal signs have a line through them by the way.

OpenStudy (pfenn1):

\[\frac{-14x^3}{x^3 - 5x^4}=\frac{-14x^3}{x^3(1-5x)}\] Can you simplify this further??

OpenStudy (anonymous):

I got D.

OpenStudy (pfenn1):

Correct.